gracesparaccio:

Jacob and Kerrie are trying to figure out how long they can live off their $84,670 savings if they spend $370 each month. They have each created an equation. Jacob’s Equation Kerrie’s Equation y = 84,670 + (370x) y = 84,670 − (370x) Which person has the correct equation to model this linear relationship? Kerrie’s equation is incorrect, because their spending will increase the amount of their savings. Kerrie’s equation is correct, because their spending will be multiplied by the number of months and then subtracted from their savings. Jacob’s equation is correct, because their spending will be multiplied by the number of months and then subtracted from their savings. Jacob’s equation is incorrect, because their spending should be a fixed amount and should not be multiplied.

Shadow:

Hello @gracesparaccio

gracesparaccio:

hi

Shadow:

Do you have an idea of how to start this question or are you not sure how to start?

gracesparaccio:

not shure on how to start

Shadow:

Let y = how much money I have Let x = the number of days Say for example, I am given an allowance of $10 dollars a week. I spend $1 a day on candy. Therefore if I am three days into the week, I have spent $3. y = 10 - x y = 10 - 3 y = 7

gracesparaccio:

y =10-3 ??

Shadow:

Yes I am illustrating how the concept works. Does that example make sense to you?

gracesparaccio:

yes thank you

Shadow:

What do you think the answer is then?

Shadow:

Or are you still unsure?

gracesparaccio:

b i think

Shadow:

You are correct.

gracesparaccio:

tyy

Shadow:

No problem (:

gracesparaccio:



gracesparaccio:

lol